OpenStudy (anonymous):

Solve the equations: x-y-5=0, x^2+2xy+y^2=9. . . . ANSWER = {(4,-1),(1,-4)}

OpenStudy (anonymous):

HELP...

jhonyy9 (jhonyy9):

opinion from this ???

OpenStudy (anonymous):

i want to solve this by substitution

jhonyy9 (jhonyy9):

there is a system of equations

jhonyy9 (jhonyy9):

so if you wann to solve it by substitution method what you need to do for ?

OpenStudy (anonymous):

how will i do this by substitution i dont understand the equations... thats the point of asking

OpenStudy (zzr0ck3r):

hint: if you have y = something, then you can plug that in for y in the other equation.

OpenStudy (anonymous):

like y=x-5 right??

jhonyy9 (jhonyy9):

not is necessary so if you know how you need ,how you can substitute the y value or x value from one equation in place of y or x inside second equation ok ?

OpenStudy (zzr0ck3r):

note if i had x+y = 3, and x+2y = 12 solve for y in the first equation y = 3-x plug this in for y in the second equation x+2(3-x) = 12 x+6 - 2x = 12 -x = 6 x = -6 now i plug this into either equation and solve for y

jhonyy9 (jhonyy9):

koli there you was right continue it

OpenStudy (anonymous):

y=x-5 ??

jhonyy9 (jhonyy9):

yes is right sure

OpenStudy (anonymous):

what should i do with this put it in second equation?

OpenStudy (zzr0ck3r):

yes

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

ok wait

jhonyy9 (jhonyy9):

just you need to make the calcule

OpenStudy (anonymous):

x^2+2x^2-10x+x^2-10x+25=0 by substituting y

OpenStudy (anonymous):

4x^2-20x+25=0 by simplifying further

jhonyy9 (jhonyy9):

yes and continue it

OpenStudy (anonymous):

(x-5)(4x-5) by factorizing right??

OpenStudy (anonymous):

x=5 and x=5/4

jhonyy9 (jhonyy9):

so if you have maked right,correct these calcules than is right sure hope so much that you have understood it all

OpenStudy (anonymous):

not yet

jhonyy9 (jhonyy9):

yes because now if you have the x values you need to calcule again the y values ok ?

OpenStudy (zzr0ck3r):

I thought your x values were 4 and 1

OpenStudy (anonymous):

not mine x values that is the book's answer

jhonyy9 (jhonyy9):

so if you know that x=y+5 so than y= ?

OpenStudy (anonymous):

it means i am doing wrong?

jhonyy9 (jhonyy9):

koli you dont wann writing here your calcules for can being checked ?

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

so if x=y+5 than the second equation will be (y+5)^2 +2(y+5)y +y^2 =9 so can you continue it ?

OpenStudy (zzr0ck3r):

he said show us your work

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

the book is wrong it seems because this will factor into c(x+b)(x+a) = 0 so both our x's will be positive

OpenStudy (zzr0ck3r):

err negative*

jhonyy9 (jhonyy9):

koli let's go please now here

OpenStudy (anonymous):

i did a mistake

OpenStudy (zzr0ck3r):

show your work to him:)

OpenStudy (anonymous):

i forgot about 9 in eqn 2 opps wait i am solving again

jhonyy9 (jhonyy9):

(y+5)^2 +2(y+5)y +y^2 =9 so can you continue it ?

jhonyy9 (jhonyy9):

y^2 +10y +25 +2y^2 +10y +y^2 =9 continue it please

OpenStudy (anonymous):

4y^2+20y+16=0

jhonyy9 (jhonyy9):

yes and continue it

jhonyy9 (jhonyy9):

come on ,please !!! this is easy

OpenStudy (anonymous):

wait i am doing

jhonyy9 (jhonyy9):

write it here please

OpenStudy (anonymous):

y=-1 and y=-4

jhonyy9 (jhonyy9):

so divide both sides by 4 and will get ???

OpenStudy (anonymous):

i did it by quadratic formula

OpenStudy (anonymous):

and i am substituting the values of y in x=y+5

jhonyy9 (jhonyy9):

4y^2+20y+16=0 /4 y^2 +5y +4 =0 factoriz it and will get (y+4)(y+1)=0 right ?

OpenStudy (anonymous):

ANSWER = {(4,-1),(1,-4)} DID IT!!

OpenStudy (anonymous):

thanks for helping me out

jhonyy9 (jhonyy9):

you have got it right understood it all ?

OpenStudy (anonymous):

there is a question a very tuff one not like this question i have already asked it before but nobody could answer it, i waited for hours and nobody helped me out i am posting the question please can u help

jhonyy9 (jhonyy9):

but firstly this equation you have understood it all right sure ?

OpenStudy (anonymous):

yeah i got it

jhonyy9 (jhonyy9):

ok so than i like help you so if i will can to do it what is your question ?